OpenStudy (anonymous):

for emlimination method 3x+2y=6 6x+4y=-3 i got 36=9 is that correct

OpenStudy (anonymous):

for elimination i seem to get 0=-15 with the following work

OpenStudy (anonymous):

Solve the first equation for x or y then insert that value for the x or y in the sceond one.

OpenStudy (anonymous):

that would be using substitution though

OpenStudy (anonymous):

these lines are parallel anyway so there is no solution

OpenStudy (anonymous):

3x+2y=6 3y = 6 - 3x y = (6-3x)/3

OpenStudy (anonymous):

insert (6-3x)/3 for y to solve for x

OpenStudy (anonymous):

@insanitycat i actually had that answer in the beginning but for some reason i changed it thanks for the help :)

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

It is 0=9 3x + 2y = 6 6x + 4y = -3 <------your problem Step One: 2(3x + 2y = 6) 6x + 4y = -3 Step Two: 6x + 4y = 12 6x + 4y = -3 Step Three: 6x - 6x = 0x 4y - 4y = 0y 12 - (-3) = 9 That's how it is done! Hope It helps!!! :)